# HoSNNs: Adversarially-Robust Homeostatic Spiking Neural Networks with Adaptive Firing Thresholds

### v3

@@ Abstract
While spiking neural networks (SNNs) offer a promising neurally-inspired model of computation, they are vulnerable to adversarial attacks. We present the first study that draws inspiration from neural homeostasis to design a threshold-adapting leaky integrate-and-fire (TA-LIF) neuron model and utilize TA-LIF neurons to construct the adversarially robust homeostatic SNNs (HoSNNs) for improved robustness. The TA-LIF model incorporates a self-stabilizing dynamic thresholding mechanism, offering a local feedback control solution to the minimization of each neuron’s membrane potential error caused by adversarial disturbance. Theoretical analysis demonstrates favorable dynamic properties of TA-LIF neurons in terms of the bounded-input bounded-output stability and suppressed time growth of membrane potential error, underscoring their superior robustness compared with the standard LIF neurons. When trained with weak FGSM attacks ($\epsilon=2/255$) and tested with much stronger PGD attacks ($\epsilon=8/255$), our HoSNNs significantly improve model accuracy on several datasets: from 30.54% to 74.91% on FashionMNIST, from 0.44% to 35.06% on SVHN, from 0.56% to 42.63% on CIFAR10, from 0.04% to 16.66% on CIFAR100, over the conventional LIF-based SNNs.
@@ 1 Introduction
While neural network models have gained widespread adoption across many domains, a glaring limitation of these models has also surfaced — vulnerability to adversarial attacks Szegedy et al. (2013); Madry et al. (2017). Subtle alterations in the input can trick a well-tuned neural network to produce misleading predictions, particularly for mission critical applications Chakraborty et al. (2018). This vulnerability is shared by both artificial neural networks (ANNs) and spiking neural networks (SNNs) Sharmin et al. (2019, 2020); Ding et al. (2022), and stands in stark contrast to the inherent robustness of biological nervous systems, prompting interesting questions: Why is the human brain immune to such adversarial noise? Can we leverage biological principles to bolster the resilience of artificial networks?
@@
Motivated by these questions, we offer a new perspective that connects adversarial robustness with homeostatic mechanisms prevalent in living organisms. Homeostasis maintains essential regulatory variables within a life-sustaining range Bernard (1865); Cooper (2008); Pennazio (2009); Jänig (2022), and is crucial for stabilizing neural activity Turrigiano and Nelson (2004), supporting neurodevelopment Marder and Goaillard (2006), and minimizing noisy information transfer Woods and Wilson (2013); Modell et al. (2015). Although some studies have investigated homeostasis in SNNs, such as the generalized leaky-integrate-and-fire (GLIF) models AllenInstitute (2018); Bellec et al. (2018, 2020); Teeter et al. (2018), no prior work has connected homeostasis with adversarial robustness.
We aim to close this gap by exploring an online biologically-plausible defense solution based on homeostasis. The proposed approach differs from common practices such as adversarial training in a major way, it explicitly builds a localized neural-level self-adapting feedback mechanism into the dynamic operation of the proposed HoSNNs. We view the time-evolving state, i.e., membrane potential $u_{i}(t|x)$ of each spiking neuron $i$ in a well trained network as its representation of the semantics of the received clean input $x$. Perturbed membrane potential $u_{i}(t|x^{\prime})$ resulting from an adversarial input $x^{\prime}$ corresponds to distorted semantics and can ripple through successive layers to mislead the network output Li et al. (2021); Rabanser et al. (2019); Nadhamuni (2021); Shu et al. (2020); Fawzi et al. (2016); Ford et al. (2019); Kang et al. (2019); Ilyas et al. (2019). For a given pair $(x,x^{\prime})$, we ensure adversarial robustness by minimizing the total induced membrane potential perturbation $E_{i}(x,x^{\prime})$.
@@
We visualize the working of the proposed HoSNNs versus standard LIF-based SNNs in image classification using several CIFAR-10 images in Figure 2. While the adversarial images generated by the Projected Gradient Descent (PGD) Madry et al. (2017) attack can completely mislead the attention of the LIF SNN, the proposed homeostasis helps the HoSNN focus on parts of the input image strongly correlated with the ground truth class label, leading to significantly improved adversarial robustness as demonstrated in detail in Section 4.
Figure 2: The heatmaps generated by Grad-CAM Selvaraju et al. (2019) highlight the regions of the input image that most significantly influence the classification decisions of a standard SNN and proposed HoSNN based on the VGG architecture for a set of CIFAR-10 images. The adversarial images are generated using PGD with an attack strength of $\epsilon=6/255$. HoSNN can still maintain attention to the target object under attack. The original Grad-CAM method is adapted for SNNs, where each heatmap is computed with respect to the input image averaged across multiple time frames.
@@ 2.1 Adversarial Attacks
Two notable adversarial attack techniques are the Fast Gradient Sign Method (FGSM) Goodfellow et al. (2014) and Projected Gradient Descent (PGD) Madry et al. (2017) method. Let $x$ be an original input, $y$ be the true label, $L(\theta,x,y)$ the loss function with network parameters $\theta$, and $\epsilon$ a small perturbation magnitude, FGSM generates a perturbed input, or an adversarial input example $x^{\prime}$ by: $x^{\prime}=x+\epsilon\cdot sign[\nabla_{x}L(\theta,x,y)]$. PGD is essentially an iterative FGSM. With $x_{n}$ as the perturbed input in the $n$-th iteration and $\alpha$ as the step size, $Proj_{x+\epsilon}\{\cdot\}$ as the projection within the $\epsilon$-ball of $x$, PGD updates the input by: $x_{n+1}=Proj_{x+\epsilon}\{x_{n}+\alpha\cdot sign[\nabla_{x}L(\theta,x_{n},y)]\}$. Other gradient-based attacks such as the RFGSM Tramèr et al. (2017), Basic Iterative Method (BIM) Kurakin et al. (2018), and DeepFool Moosavi-Dezfooli et al. (2016) work in a similar fashion. Beyond gradient-based methods Biggio and Roli (2018), a significant area of concern is black-box attacks.
@@ 2.2 Defense Methods
Adversarial training is one of the most widely adopted defense methods Madry et al. (2017), which retrains a model by using a mixture of clean and adversarial examples. Randomization Xie et al. (2017) introduces stochasticity during inference and can circumvent precise adversarial attacks. The projection technique of Mustafa et al. (2019) reverts adversarial attacks back to a safer set. Lastly, one may first detect the presence of adversarial attack and subsequently cope with it Metzen et al. (2017). However, these defense methods are not free of limitations Akhtar and Mian (2018). Adversarial training relies on precise gradient information and is not biologically plausible Lillicrap et al. (2020). Randomization, projection, and detection strategies do not fundamentally address the inherent vulnerabilities of neural networks.
@@ 2.3 Spiking Neuron Models
SNNs allow for spike-based communication and computationFurber et al. (2014); Gerstner and Kistler (2002); Deng et al. (2020) and often leverage the Leaky Integrate-and-Fire (LIF) model for each neuron $i$ with the membrane time constant $\tau_{m}$:
@@
$s_{i}(t)=\begin{cases}+\infty&\text{if }u_{i}(t)\geq V^{i}_{th}(t),\\ 0&\text{otherwise}\end{cases}=\sum_{f}\delta(t-t_{i}^{f})$ (2)
where $\delta(\cdot)$ is the Dirac function, and $t_{i}^{f}$ is a postsynaptic spike time. The LIF model uses a constant firing threshold. Generalized LIF (GLIF) models employ a tunable threshold with short-term memory, which increases with every emitted output spike, and subsequently decays exponentially back to the baseline threshold AllenInstitute (2018); Bellec et al. (2018, 2020); Teeter et al. (2018). However, these models neither consider adversarial robustness nor provides mechanisms to discern “abnormal" neural activity from “normal" activity.
@@ 2.4 Spiking Neural Networks Robustness
While there’s been growing interest in spiking neural networks Imam and Cleland (2020); Pei et al. (2019), empirical studies have demonstrated that SNNs exhibit similar susceptibilities to adversarial attacks Sharmin et al. (2019); Ding et al. (2022). A line of research has explored porting defensive strategies developed for ANNs to SNNs. To improve the robustness of SNNs, Kundu et al. (2021) proposed a SNN training algorithm jointly optimizing firing thresholds and weights, and Liang et al. (2022) proposed certified training. Ding et al. (2022) enhanced adversarial training by using a Lipschitz constant regularizer. However, these methods have not fully addressed the challenges of ensuring adversarial robustness. Additionally, they are computationally expensive and lack biological plausibility. Another research direction has focused on studying the inherent robustness of SNNs not seen in their ANN counterparts, and factors impacting robustness. Sharmin et al. (2020) recognized the inherent resistance of SNNs to gradient-based adversarial attacks. El-Allami et al. (2021) investigated the impact of key network parameters such as firing voltage thresholds on robustness. Chowdhury et al. (2021) demonstrated the LIF model’s noise-filtering capability. Li et al. (2022) explored network inter-layer sparsity, and Xu et al. (2022) examined the effects of surrogate gradient techniques on white-box attacks. While these studies have shed light on aspects of SNNs relevant to robustness, effective defense strategies are yet to be developed.
@@ 3.1 Adversarial Robustness as a Membrane Potential Error Minimization Problem
In a well trained network, we view the time-evolving state, i.e., membrane potential $u_{i}(t|x)$ of each spiking neuron $i$, over $T$ timesteps as its representation of the semantics of the received clean input $x$. An adversarial input $x^{\prime}=x+\epsilon\delta x$, where $\epsilon$ is the attack budget (strength), and $\delta x$ is a carefully crafted adversarial noise, may lead to a perturbed membrane potential $u_{i}(t|x^{\prime})$, which corresponds to distorted semantics and can ripple through successive layers to mislead the network’s decision Li et al. (2021); Rabanser et al. (2019); Nadhamuni (2021); Shu et al. (2020); Fawzi et al. (2016); Ford et al. (2019); Kang et al. (2019); Ilyas et al. (2019).
@@
$\min E_{i}(x,x^{\prime})=\sum_{i=0}^{N}\sum_{t=0}^{T}\frac{1}{2}\parallel u_{i}(t|x^{\prime})-u_{i}(t|x)\parallel^{2}$ (3)
However, there exist two challenges in formulating and solving (3). Firstly, since the model is oblivious about the attack, it is impossible to determine how the adversarial input $x^{\prime}$ is generated, whether it has a corresponding clean input $x$, and what $x$ is if it exists. As such, $u_{i}(t|x)$ is unknown, which serves as a clean reference in (3). We address this problem by inducing the notion of Neural Dynamic Signature (NDS), the neuron’s membrane potential averaged over a given clean training dataset $\mathcal{D}$, to provide an anchor for stabilizing membrane potential. Secondly, since externally generated attacks are not known a priori, it is desirable to suppress the perturbation of each neuron’s membrane potential in an online manner. For this, we propose a new type of spiking neurons, called threshold-adapting leaky integrate-and-fire (TA-LIF) neurons with a properly designed firing threshold voltage dynamics that serves as a homeostatic control to suppress undesirable membrane potential perturbations. We discuss these two techniques next.
@@ 3.2 Neural Dynamic Signature (NDS) as an Anchor
We also define the network-level NDS, $\mathcal{U}_{\text{NET}}(\mathcal{D})\triangleq\{\bm{u}_{i}(\mathcal{D})\}_{i=1}^{N}$, as the collection of NDS vectors in a well-trained LIF-based SNN comprising $N$ neurons. A densely populated SNN with $N$ neurons typically has $O(N^{2})$ weight parameters. In contrast, $\mathcal{U}_{\text{NET}}(\mathcal{D})$ scales as $O(NT)$. Recent algorithms have enabled training of high-accuracy SNNs with short latency operating over a small number of time steps, e.g., 5 to 10 Zhang and Li (2020). Consequently, the storage overhead of NDS remains manageable.
@@
While serving as an anchor signal, the dynamics of NDS provides a basis for understanding the property of the proposed TA-LIF neurons. As NDS $u^{*}_{i}(t|\mathcal{D})$ of neuron $i$ is derived from the well-trained LIF SNN, we take expectation of the LIF dynamic equation (1) with a static firing threshold $V_{th}$ across the entire training distribution $\mathcal{D}$ while simplifying the dynamics by approximating the effects of firing:

### v4

@@ Abstract
While spiking neural networks (SNNs) offer a promising neurally-inspired model of computation, they are vulnerable to adversarial attacks. We present the first study that draws inspiration from neural homeostasis to design a threshold-adapting leaky integrate-and-fire (TA-LIF) neuron model and utilize TA-LIF neurons to construct the adversarially robust homeostatic SNNs (HoSNNs) for improved robustness. The TA-LIF model incorporates a self-stabilizing dynamic thresholding mechanism, offering a local feedback control solution to the minimization of each neuron’s membrane potential error caused by adversarial disturbance. Theoretical analysis demonstrates favorable dynamic properties of TA-LIF neurons in terms of the bounded-input bounded-output stability and suppressed time growth of membrane potential error, underscoring their superior robustness compared with the standard LIF neurons. When trained with weak FGSM attacks ($\epsilon=2/255$), our HoSNNs significantly outperform conventionally trained LIF-based SNNs across multiple datasets. Furthermore, under significantly stronger PGD7 attacks ($\epsilon=8/255$), HoSNN achieves notable improvements in accuracy, increasing from 30.90% to 74.91% on FashionMNIST, 0.44% to 36.82% on SVHN, 0.54% to 43.33% on CIFAR10, and 0.04% to 16.66% on CIFAR100.
@@ 1 Introduction
While neural network models have gained widespread adoption across many domains, a glaring limitation of these models has also surfaced — vulnerability to adversarial attacks (Szegedy et al., 2013; Madry et al., 2017). Subtle alterations in the input can trick a well-tuned neural network into producing misleading predictions, particularly for mission-critical applications (Chakraborty et al., 2018). This vulnerability is shared by both artificial neural networks (ANNs) and spiking neural networks (SNNs) (Sharmin et al., 2019; 2020; Ding et al., 2022), and stands in stark contrast to the inherent robustness of biological nervous systems, prompting interesting questions: Why is the human brain immune to such adversarial noise? Can we leverage biological principles to bolster the resilience of artificial networks?
@@
Motivated by these questions, we offer a new perspective that connects adversarial robustness with homeostatic mechanisms prevalent in living organisms. Homeostasis maintains essential regulatory variables within a life-sustaining range (Bernard, 1865; Cooper, 2008; Pennazio, 2009; Jänig, 2022), and is crucial for stabilizing neural activity (Turrigiano & Nelson, 2004), supporting neurodevelopment (Marder & Goaillard, 2006), and minimizing noisy information transfer (Woods & Wilson, 2013; Modell et al., 2015). Although some studies have investigated homeostasis in SNNs, such as the generalized leaky-integrate-and-fire (GLIF) models (AllenInstitute, 2018; Bellec et al., 2018; 2020; Teeter et al., 2018), no prior work has connected homeostasis with adversarial robustness.
We aim to close this gap by exploring an online biologically-plausible defense solution based on homeostasis. The proposed approach differs from common practices such as adversarial training in a major way, it explicitly builds a localized neural-level self-adapting feedback mechanism into the dynamic operation of the proposed HoSNNs. We view the time-evolving state, i.e., membrane potential $u_{i}(t|x)$ of each spiking neuron $i$ in a well trained network as its representation of the semantics of the received clean input $x$. Perturbed membrane potential $u_{i}(t|x^{\prime})$ resulting from an adversarial input $x^{\prime}$ corresponds to distorted semantics and can ripple through successive layers to mislead the network output (Li et al., 2021; Rabanser et al., 2019; Nadhamuni, 2021; Shu et al., 2020; Fawzi et al., 2016; Ford et al., 2019; Kang et al., 2019; Ilyas et al., 2019a). For a given pair $(x,x^{\prime})$, we ensure adversarial robustness by minimizing the total induced membrane potential perturbation $E_{i}(x,x^{\prime})$.
@@
We visualize the working of the proposed HoSNNs versus standard LIF-based SNNs in image classification using several CIFAR-10 images in Figure 2. While the adversarial images generated by the Projected Gradient Descent (PGD) (Madry et al., 2017) attack can completely mislead the attention of the LIF SNN, the proposed homeostasis helps the HoSNN focus on parts of the input image strongly correlated with the ground truth class label, leading to significantly improved adversarial robustness as described later.
Figure 2: The heatmaps generated by Grad-CAM (Selvaraju et al., 2019) highlight the regions of the input image that most significantly influence the classification decisions of a standard SNN and proposed HoSNN based on the VGG architecture for a set of CIFAR-10 images. The adversarial images are generated using PGD7 with strength of $\epsilon=6/255$. HoSNN can still maintain attention to the target object under attack.
@@ 2.1 Adversarial Attacks
Two notable adversarial attack techniques are the Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2014) and Projected Gradient Descent (PGD) (Madry et al., 2017) method. Let $x$ be an original input, $y$ be the true label, $L(\theta,x,y)$ the loss function with network parameters $\theta$, and $\epsilon$ a small perturbation magnitude, FGSM generates a perturbed input, or an adversarial input example $x^{\prime}$ by: $x^{\prime}=x+\epsilon\cdot sign[\nabla_{x}L(\theta,x,y)]$. PGD is essentially an iterative FGSM. With $x_{n}$ as the perturbed input in the $n$-th iteration and $\alpha$ as the step size, $Proj_{x+\epsilon}\{\cdot\}$ as the projection within the $\epsilon$-ball of $x$, PGD updates the input by: $x_{n+1}=Proj_{x+\epsilon}\{x_{n}+\alpha\cdot sign[\nabla_{x}L(\theta,x_{n},y)]\}$. Other gradient-based attacks such as the RFGSM (Tramèr et al., 2017), Basic Iterative Method (BIM) (Kurakin et al., 2018), and DeepFool (Moosavi-Dezfooli et al., 2016) work in a similar fashion, APGD (Croce & Hein, 2020b). Beyond gradient-based methods, a significant area of concern is black-box attacks (Biggio & Roli, 2018).
Some recent studies focused on delivering more powerful attacks in SNNs. Bu et al. (2023) proposed RGA, which leveraged rate-coding in LIF neurons to generate more effective adversarial attacks with a time-extended enhancement. Hao et al. (2024) proposed HART, which leveraged a hybrid gradient calculation that simultaneously incorporates rate-based gradients and timing-based temporal gradients on SNNs.
@@ 2.2 Defense Methods
Adversarial training is one of the most widely adopted defense methods (Madry et al., 2017), which retrains a model by using a mixture of clean and adversarial examples. Randomization (Xie et al., 2017) introduces stochasticity during inference and can circumvent precise adversarial attacks. The projection technique of (Mustafa et al., 2019) reverts adversarial attacks back to a safer set. Lastly, one may first detect the presence of adversarial attack and subsequently cope with it (Metzen et al., 2017a). However, these defense methods are not free of limitations (Akhtar & Mian, 2018). Adversarial training relies on precise gradient information and is not biologically plausible (Lillicrap et al., 2020). Randomization, projection, and detection strategies do not fundamentally address the inherent vulnerabilities of ANNs.
@@ 2.3 Spiking Neuron Models
SNNs allow for spike-based communication and computation (Furber et al., 2014; Gerstner & Kistler, 2002; Deng et al., 2020) and often leverage the Leaky Integrate-and-Fire (LIF) model for each neuron $i$ with the membrane time constant $\tau_{m}$:
@@
$s_{i}(t)=\begin{cases}+\infty&\text{if }u_{i}(t)\geq V^{i}_{th}(t)\\ 0&\text{otherwise}\end{cases}=\sum_{f}\delta(t-t_{i}^{f})$ (2)
where $\delta(\cdot)$ is the Dirac function, and $t_{i}^{f}$ is a postsynaptic spike time. With $\tau_{s}$ denoting the synaptic time constant, the evolution of the generated postsynaptic current (PSC) $a_{j}(t)$ is described by:
$\tau_{s}\frac{da_{j}(t)}{dt}=-a_{j}(t)+s_{j}(t)$ (3)
The LIF model uses a constant firing threshold. Generalized LIF (GLIF) models employ a tunable threshold with short-term memory, which increases with every emitted output spike, and subsequently decays exponentially back to the baseline threshold (AllenInstitute, 2018; Bellec et al., 2018; 2020; Teeter et al., 2018). However, these models neither consider adversarial robustness nor provide mechanisms to discern "abnormal" from "normal" neural activity.
@@ 2.4 Spiking Neural Networks Robustness
While there’s been growing interest in spiking neural networks (Imam & Cleland, 2020; Pei et al., 2019), empirical studies have demonstrated that SNNs exhibit similar susceptibilities to adversarial attacks (Sharmin et al., 2019; Ding et al., 2022). A line of research has explored porting defensive strategies developed for ANNs to SNNs. To improve the robustness of SNNs, Kundu et al. (2021) proposed a SNN training algorithm jointly optimizing firing thresholds and weights, and Liang et al. (2022) proposed certified training. Ding et al. (2022) enhanced adversarial training by using a Lipschitz constant regularizer. Özdenizci & Legenstein (2023) introduces an adversarially robust ANN-to-SNN conversion algorithm that initializes the SNN with adversarially pre-trained ANN weights, followed by robust fine-tuning. Liu et al. (2024) improved SNN’s adversarial robustness by adding a gradient sparsity regularization term in the loss function. However, these methods have not fully addressed the challenges of ensuring adversarial robustness. Additionally, they are computationally expensive and lack biological plausibility.
Another research direction has focused on studying the inherent robustness of SNNs not seen in their ANN counterparts, and factors impacting robustness. Sharmin et al. (2020) recognized the inherent resistance of SNNs to gradient-based adversarial attacks. El-Allami et al. (2021) investigated the impact of key network parameters such as firing voltage thresholds on robustness. Chowdhury et al. (2021) demonstrated the LIF model’s noise-filtering capability. Li et al. (2022) explored network inter-layer sparsity. Xu et al. (2022) examined the effects of surrogate gradient techniques on white-box attacks. While these studies have shed light on aspects of SNNs relevant to robustness, effective defense strategies are yet to be developed.
Similar to ours, some studies improve robustness from stability and biological and rational perspectives. Ding et al. (2024a) enhanced adversarial robustness by reducing the mean square of perturbations in the last neuron layer. Inspired by Stochastic Gating Mechanisms, Ding et al. (2024b) introduced randomness into spike transmission by simulating the probabilistic opening and closing of synaptic and ion channel gates. The fundamental differences between our work and theirs are: 1.We adjust the threshold in a passive and online manner instead of deliberately adding additional optimization loss terms to the loss function as (Ding et al., 2024a), which significantly reduces the additional calculations for training and inference. 2.Our method does not introduce additional randomness as (Ding et al., 2024b). Thus, our method can retain high clean accuracy and ensures that the robustness does not come from gradient obfuscation (Athalye et al., 2018).
@@ 3.1 Adversarial Robustness as a Membrane Potential Error Minimization Problem
In a well-trained network, we view the time-evolving state, i.e., membrane potential $u_{i}(t|x)$ of each spiking neuron $i$, over $T$ timesteps as its representation of the semantics of the received clean input $x$. An adversarial input $x^{\prime}=x+\delta x,\;s.t.\;\delta x<\epsilon$, where $\epsilon$ is the attack budget (strength), and $\delta x$ is a carefully crafted adversarial noise, may lead to a perturbed membrane potential $u_{i}(t|x^{\prime})$, which corresponds to distorted semantics and can ripple through successive layers to mislead the network’s decision (Li et al., 2021; Rabanser et al., 2019; Nadhamuni, 2021; Shu et al., 2020; Fawzi et al., 2016; Ford et al., 2019; Kang et al., 2019; Ilyas et al., 2019a).
@@
$\min E_{i}(x,x^{\prime})=\sum_{i=0}^{N}\sum_{t=0}^{T}\parallel u_{i}(t|x^{\prime})-u_{i}(t|x)\parallel^{2}$ (4)
However, there exist two challenges in formulating and solving equation 4. Firstly, since the model is oblivious about the attack, it is impossible to determine how the adversarial input $x^{\prime}$ is generated, whether it has a corresponding clean input $x$, and what $x$ is if it exists. As such, $u_{i}(t|x)$ is unknown, which serves as a clean reference in equation 4. We address this problem by inducing the notion of Neural Dynamic Signature (NDS), the neuron’s membrane potential averaged over a given clean training dataset $\mathcal{D}$, to provide an anchor for stabilizing membrane potential. Secondly, since externally generated attacks are not known a priori, it is desirable to suppress the perturbation of each neuron’s membrane potential in an online manner. For this, we propose a new type of spiking neurons, called threshold-adapting leaky integrate-and-fire (TA-LIF) neurons with a properly designed firing threshold voltage dynamics that serves as a homeostatic control to suppress undesirable membrane potential perturbations. We discuss these two techniques next.
@@ 3.2 Neural Dynamic Signature (NDS) as an Anchor
Eq 4 describes an "ideal" optimization problem. When $E_{i}(x,x^{\prime})=0$, all neuron activities under adversarial sample input are the same as normal sample input, this can certainly achieve "adversarial robustness" in theory, but it is impossible to achieve in practice. There are two main limitations:(1) For a trained network, the intensity and type of attack are determined by external attackers, which means that x’ has a huge range of variation. (2) For a trained network, the attack sample x cannot be determined in advance, which means that it is impossible to obtain an accurate $u_{i}(t|x)$. Under the constraints of these two problems, we still hope to adopt the idea of Eq 4. A feasible approximation is that we use the available average value of the training data set $\mathbb{E}_{x\sim\mathcal{D}}\left[u_{i}(t|x)\right]$ as an approximation of the unavailable $u_{i}(t|x)$.
@@
We also define the network-level NDS, $\mathcal{U}_{\text{NET}}(\mathcal{D})\triangleq\{\bm{u}_{i}(\mathcal{D})\}_{i=1}^{N}$, as the collection of NDS vectors in a well-trained LIF-based SNN comprising $N$ neurons. A densely populated SNN with $N$ neurons typically has $O(N^{2})$ weight parameters. In contrast, $\mathcal{U}_{\text{NET}}(\mathcal{D})$ scales as $O(NT)$. Recent algorithms have enabled training of high-accuracy SNNs with short latency operating over a small number of time steps, e.g., 5 to 10 (Zhang & Li, 2020). Consequently, the storage overhead of NDS remains manageable.
@@
While serving as an anchor signal, the dynamics of NDS provides a basis for understanding the property of the proposed TA-LIF neurons. As NDS $u^{*}_{i}(t|\mathcal{D})$ of neuron $i$ is derived from the well-trained LIF SNN, we take expectation of the LIF dynamic equation 1 with a static firing threshold $V_{th}$ across the entire training distribution $\mathcal{D}$ while simplifying the dynamics by approximating the effects of firing:
@@
Hypothesis of NDS
We hypothesize that in SNNs, the original neural activity within the training set holds reference significance for mitigating adversarial perturbations. This idea finds support in certain works within the Artificial Neural Network (ANN) domain, which suggest that detecting and modifying neuron activation values in feature space can alleviate adversarial issues (Silva & Najafirad, 2020; Metzen et al., 2017b; Zhang et al., 2020). Some studies deny this view (Carlini & Wagner, 2017; Tramèr, 2022; Ilyas et al., 2019b; Li et al., 2021). In SNNs, whether membrane potential sequences can effectively address adversarial attack problems remains unclear and needs further investigation.
Here we provide an intuitive understanding of NDS’s effectiveness. Consider an adversarial sample $x^{\prime}=x+\delta x$, where the perturbation $\delta x$ represents the adversarial attack. As the strength of the perturbation diminishes ($\delta x\rightarrow 0$), the sample gradually converges to the original clean sample ($x^{\prime}\rightarrow x$). Consequently, the neural activity of the attacked network, denoted as $f(x^{\prime})$, will approach the clean neural activity, $f(x)$, i.e., $f(x^{\prime})\rightarrow f(x)$. This alignment between the attacked and clean neural activity suggests that the network’s classification result will likewise revert to the correct prediction, $y^{\prime}\rightarrow y$. Therefore, if we can identify a method to reduce the discrepancy between the neural activity of the attacked network and that of the clean network, we may be able to mitigate the impact of adversarial attacks on the classification outcome.